# Non-thermal emission from mildly relativistic tidal ejecta of compact objects merger

### v1

@@ 5.3 Sky image
To further validate our semi-analytic calculation, in Fig. 4, we show an example of the flux emitted from different annuli obtained numerically and semi-analytically for spherical ejecta (we considered $\theta_{\text{open}}=90^{\circ}$), along with the analytic estimate of the image radius from (Sadeh et al., 2024). Due to relativistic beaming and time arrival effects, the image is a relatively narrow ring.
Figure 4: The fraction of flux that is emitted at different times by different annuli, $df_{\nu}/dr$, for $p=2.2,s_{\text{ft}}=7,\gamma_{0}=1.15$. In solid lines: full numerical 1D calculation. In dashed lines: our semi-analytic scheme. $r$, the radius of the annulus (its transverse distance to the line of sight) is normalized to $R_{R}=ct_{R}$ (approximately the radius at which the post-shock plasma momentum drops to $\gamma\beta=1$), where $t_{R}\equiv\left(\frac{M_{R}}{16\pi nm_{p}c^{3}}\right)^{\frac{1}{3}}$. The $+$’s are the analytic estimate of the image radius, given in Sadeh et al. (2024).
@@ 6 Results
In Figs. 5-6, we provide several examples of the obtained radio maps and the corresponding light curves from the semi-analytic calculation for several values of $\theta_{\text{v}}$ and typical parameters expected for NSBH merger, $\{s_{\text{ft}},s_{\text{KN}},\beta_{0},n,M_{R},\varepsilon_{e},\varepsilon_{B},\theta_{\text{open}}\}=\{7,1.5,0.3,10^{-2}\text{cm}^{-3},10^{-4}M_{\odot},10^{-1},10^{-2},20^{\circ}\}$.
Figure 5: The observed sky radio map for a distance of $100$Mpc, frequency of $1$GHz at $t=7110$days, for parameters expected for NSBH merger, $\{s_{\text{ft}},s_{\text{KN}},\beta_{0},n,M_{0},\varepsilon_{e},\varepsilon_{B},p,\theta_{\text{open}}\}=\{7,1.5,0.3,10^{-2}\text{cm}^{-3},8\times 10^{-2}M_{\odot},10^{-1},10^{-2},2.2,20^{\circ}\}$, and for different observation angles $\theta_{\text{v}}=0^{\circ},30^{\circ},60^{\circ},90^{\circ}$, calculated by the semi-analytic method described in §3. In blue: the analytic estimation from Sadeh et al. (2024) for observed image radius (the image appears as a narrow ring in the spherical case).
@@ 7 Conclusions
In this work, we study the non-thermal emission of a toroidal (see Fig. 1), mildly-relativistic ejecta with broken power-law mass distribution, given by Eq. (1). We provided a semi-analytic calculation method for the synchrotron emission, assuming emission only from the initial opening angle of the ejecta (§3) at times in which the reverse shock propagates through the steep, fast tail. Furthermore, we provided analytic estimations for the peak time and flux in Eqs. (6) and Eq. (7). Then, by calculating the full 2D hydrodynamics (§4), we validated our semi-analytic approach in §5 for a wide range of ejecta parameter values characteristic of merger calculation results; an equatorial ejecta opening angle of $15^{\circ}<\theta<30^{\circ}$, a steep ($s_{\rm ft}>5$, $\gamma_{0}\beta_{0}\geq 0.3$) mass distribution for mildly relativistic velocities and a moderate mass distribution for the bulk of the ejecta ($31<s_{\rm KN}<$ 3, $0.1<\gamma_{0}\beta_{0}\leq 0.3$). We also used various values of electrons power-law distribution ($2.42.1<p<$ 2.4). This is the first semi-analytic method supported by full 2D relativistic numeric calculations for the expected late non-thermal emission from mergers of compact objects with a substantial ($q>1.5$) mass ratio. We found that the assumption that the freshly shocked material within the initial opening angle of the ejecta has a dominant contribution to the internal energy during the phase in which the reverse shock propagates through the fast tail is in agreement with the full 2D numerical calculation (Fig. 2). Additionally, we found that, in agreement with Eq. (7), the peak flux for toroidal ejecta is smaller by a factor of $\sim 2-3$ (varies for different viewing angles) than spherical ejecta with the same isotropic equivalent mass, while the peak time estimation from (Eq. (6)) for spherical ejecta (with the same isotropic equivalent mass) is within $20\%$ deviation from the peak time results of the full numerical calculation for toroidal ejecta presented in this paper (Fig. 3). The radio map is significantly different between the toroidal and the spherical case in which the image is a symmetric ring (Fig. 5), for $\theta_{\text{v}}\geq 20^{\circ}$ the width of the observed image (for $\theta_{\text{v}}=90^{\circ}$ the width is parallel to the equatorial plane) is comparable with the ring diameter in the spherical case (Sadeh et al., 2024), while the length (perpendicular to the width) is considerably smaller, providing a clear indication of equatorial ejecta. The different viewing angles affect the observed image orientation (see Fig. 5), providing the possibility to constrain the viewing angle in future radio map observations. For $0^{\circ}\leq\theta_{\text{v}}\leq 20^{\circ}$, the image is observed as a narrow ring (Fig. 5), similar to the spherical case. The semi-analytic scheme we presented in this work can be used to fit model parameters to future observations of light curves and radio maps by an iterative process that scans the possible parameter space, e.g., Markov chain Monte Carlo.
@@
Different mass ratios between the compact objects correspond to different ejecta geometries (Bernuzzi et al., 2020). Consequently, bounds over the ejecta geometry can constrain this mass ratio. For example, in GW170817, the estimation for the mass ratio varies between $2.61<q<$ 2.6 (for low spin assumption, it is $1.41<q<$ 1.4), while the estimation for the viewing angle from gravitational waves is $32^{\circ}\pm 10^{\circ}$. Observing the radio map of the late non-thermal emission, in this case, can potentially resolve the large uncertainty in the mass ratio.
@@ Appendix A Semi-analytic calculation description
As the ejected material expands, it initiates a forward shock that propagates into the ISM, while simultaneously, a reverse shock is generated within the ejected material itself (see Fig. 7). The deceleration imposed by the reverse shock results in a reduction of kinetic energy in the ejecta and this energy is subsequently transmitted to the heated ionized plasma within the shocked ISM.
Figure 7: A schematic illustration showing the four regions of the forward-reverse shock structure described in appendix A: (1) un-shocked ISM, (2) shocked ISM, (3) shocked ejecta, (4) un-shocked ejecta. The pressure in the unshocked regions is negligible compared to that in the shocked regions, $p_{1},p_{4}\ll p_{2},p_{3}$. The pressure and velocity are approximated in our analytic calculations as uniform within the shocked region, and the density is approximated as uniform within the two regions separated by the contact discontinuity.

### v2

@@ 5.3 Sky image
To further validate our semi-analytic calculation, in Fig. 5, we show an example of the flux emitted from different annuli obtained numerically and semi-analytically for spherical ejecta (we considered $\theta_{\text{open}}=90^{\circ}$), along with the analytic estimate of the image radius from (Sadeh & Waxman, 2024). Due to relativistic beaming and time arrival effects, the image is a relatively narrow ring. This ring radius increases with time since the blast wave expands and the relativistic beaming effect decreases. The emission from large viewing angles, which was initially negligible due to the relativistic beaming, becomes brighter as the velocity declines.
Figure 5: The fraction of flux that is emitted at different times by different annuli, $df_{\nu}/dr$, for $p=2.2,s_{\text{ft}}=7,\gamma_{0}=1.15$. In solid lines: full numerical 1D calculation. In dashed lines: our semi-analytic scheme. $r$, the radius of the annulus (its transverse distance to the line of sight) is normalized to $R_{R}=ct_{R}$ (approximately the radius at which the post-shock plasma momentum drops to $\gamma\beta=1$), where $t_{R}\equiv\left(\frac{M_{R}}{16\pi nm_{p}c^{3}}\right)^{\frac{1}{3}}$. The $+$’s are the analytic estimate of the image radius, from Sadeh & Waxman (2024).
@@ 6 Results
The mass ejection in NSBH mergers simulations typically reaches $\sim 0.1M_{\odot}$ in disruptive merger systems (Foucart et al., 2017; Kyutoku et al., 2021; Chen et al., 2024), implying an isotropic equivalent mass of $M\sim 0.3M_{\odot}$ (for a typical opening angle of $\theta_{\text{open}}=20^{\circ}$). This is consistent with values of $M_{R}=10^{-4}M_{\odot},s_{\text{ft}}=7$. In Figs. 6-7, we provide several examples of the obtained radio maps and the corresponding light curves from the semi-analytic calculation for several values of $\theta_{\text{v}}$ and typical parameters expected for NSBH merger, $\{s_{\text{ft}},s_{\text{KN}},\beta_{0},n,M_{R},\varepsilon_{e},\varepsilon_{B},\theta_{\text{open}}\}=\{7,1.5,0.3,10^{-2}\text{cm}^{-3},10^{-4}M_{\odot},10^{-1},10^{-2},20^{\circ}\}$. Notice that at the peak, the shocked plasma velocity is $\sim\beta_{0}=0.3$, and the radiation emitted from such plasma is beamed into an angle of $\sim\sin^{-1}(\frac{1}{\gamma_{0}})\approx 70^{\circ}$. This is the reason for the small difference at the peak flux between viewing angles of $0^{\circ}-30^{\circ}$, in which a larger fraction of the radiation is beamed away, to viewing angles of $60^{\circ}-90^{\circ}$. We add our analytic estimation for the light curve and image radius for a spherical ejecta case with the same equivalent isotropic mass for comparison.
Figure 6: The observed sky radio map for a distance of $100$Mpc, frequency of $1$GHz at $t=7110$days, for parameters expected for NSBH merger, $\{s_{\text{ft}},s_{\text{KN}},\beta_{0},n,M_{0},\varepsilon_{e},\varepsilon_{B},p,\theta_{\text{open}}\}=\{7,1.5,0.3,10^{-2}\text{cm}^{-3},8\times 10^{-2}M_{\odot},10^{-1},10^{-2},2.2,20^{\circ}\}$, and for different observation angles $\theta_{\text{v}}=0^{\circ},30^{\circ},60^{\circ},90^{\circ}$, calculated by the semi-analytic method described in §3. At different times, the radius of the ring changes, but the physical picture remains. In blue: the analytic estimation from Sadeh & Waxman (2024) for observed image radius (the image appears as a narrow ring in the spherical case).
@@
We find that the toroidal ejecta case is less bright by a factor of $\sim 2-3$ than the isotropic equivalent case (Fig. 7), and expected to peak on a time scale of ten years. Furthermore, the peak flux’s dependence on the viewing angle is minimal for this toroidal ejecta. Oppositely, The radio map is significantly different between the different viewing angles (Fig. 6), providing an opportunity to constrain the viewing angle in future radio map observations.
@@ 7 Conclusions
In this work, we study the non-thermal emission of a toroidal (see Fig. 1), mildly-relativistic ejecta with broken power-law mass distribution, given by Eq. (1). We provided a semi-analytic calculation method for the synchrotron emission, assuming emission only from the initial opening angle of the ejecta (§3) at times in which the reverse shock propagates through the steep, fast tail. Furthermore, we provided analytic estimations for the peak time and flux in Eqs. (6) and Eq. (7). Then, by calculating the full 2D hydrodynamics (§4), we validated our semi-analytic approach in §5 for a wide range of ejecta parameter values characteristic of merger calculation results; an equatorial ejecta opening angle of $15^{\circ}<\theta<30^{\circ}$, a steep ($s_{\rm ft}>5$, $\gamma_{0}\beta_{0}\geq 0.3$) mass distribution for mildly relativistic velocities and a moderate mass distribution for the bulk of the ejecta ($31<s_{\rm KN}<$ 3, $0.1<\gamma_{0}\beta_{0}\leq 0.3$). We also used various values of electrons power-law index, $2.42.1<p<$ 2.4, of the electron energy distribution. This is the first semi-analytic method supported by full 2D relativistic numeric calculations for the expected late non-thermal emission from mergers of compact objects with a substantial ($q>1.5$) mass ratio. We found that the full 2D numerical calculation is consistent with our assumption (Fig. 3), i.e., that during reverse shock propagation through the fast tail, the freshly shocked material within the initial opening angle of the ejecta has a dominant contribution to the total internal energy. Additionally, we found that, in agreement with Eq. (7), the peak flux for toroidal ejecta is smaller by a factor of $\sim 2-3$ (varies for different viewing angles) than spherical ejecta with the same isotropic equivalent mass, while the peak time estimation from (Eq. (6)) for spherical ejecta (with the same isotropic equivalent mass) is within $20\%$ deviation from the peak time results of the full numerical calculation for toroidal ejecta presented in this paper (Fig. 4). The radio map is significantly different between the toroidal and the spherical case in which the image is a symmetric ring (Fig. 6). For $\theta_{\text{v}}\geq 20^{\circ}$ the width of the observed image (for $\theta_{\text{v}}=90^{\circ}$ the width is parallel to the equatorial plane) is comparable with the ring diameter in the spherical case (Sadeh & Waxman, 2024). Concurrently, the length (perpendicular to the width) is considerably smaller, providing a clear indication of equatorial ejecta. The different viewing angles affect the observed image orientation (see Fig. 6), providing the possibility to constrain the viewing angle in future radio map observations. For $0^{\circ}\leq\theta_{\text{v}}\leq 20^{\circ}$, the image is observed as a narrow ring (Fig. 6), similar to the spherical case. The semi-analytic scheme we presented in this work can be used to fit all of the model parameters $\{s_{\text{ft}},\gamma_{0},n,M_{R},\varepsilon_{e},\varepsilon_{B},p,\theta_{\text{open}},\theta_{\text{v}}\}$ to future observations of light curves and radio maps by an iterative process that scans the possible parameter space, e.g., Markov chain Monte Carlo. Degeneracies between the different parameters are expected to arise during the fitting process, for example, see Eq. (6)-(7). There are two distinctions: the radio map shape is expected to provide stringent bounds over $\theta_{\text{open}}$ and $\theta_{\text{v}}$, while the synchrotron spectrum provides $p$.
@@
Different mass ratios between the compact objects correspond to different ejecta geometries (Bernuzzi et al., 2020; Bernuzzi et al., 2024). Consequently, bounds over the ejecta geometry can constrain this mass ratio. For example, in GW170817, the estimation for the mass ratio varies between $2.61<q<$ 2.6 (for low spin assumption, it is $1.41<q<$ 1.4), while the estimation for the viewing angle from gravitational waves is $32^{\circ}\pm 10^{\circ}$. Observing the radio map of the late non-thermal emission, in this case, can potentially resolve the large uncertainty in the mass ratio.
@@ Appendix A Semi-analytic calculation description
As the ejected material expands, it initiates a forward shock that propagates into the ISM, while simultaneously, a reverse shock is generated within the ejected material itself (see Fig. 8). The deceleration imposed by the reverse shock reduces kinetic energy in the ejecta, which is subsequently transmitted to the heated ionized plasma within the shocked ISM.
Figure 8: A schematic illustration showing the four regions of the forward-reverse shock structure described in Appendix
A: (1) un-shocked ISM, (2) shocked ISM, (3) shocked ejecta, (4) un-shocked ejecta. The pressure in the unshocked regions is negligible compared to that in the shocked regions, $p_{1},p_{4}\ll p_{2},p_{3}$. The pressure and velocity are approximated in our analytic calculations as uniform within the shocked region, and the density is approximated as uniform within the two regions separated by the contact discontinuity.